OpenStudy (anonymous):

A pair of equations is shown below: y = 3x − 5 y = 6x − 8 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Help explain how to do this..

Mehek (mehek14):

substitution put 3x- 5 for y in the second equation

Mehek (mehek14):

so it becomes \(3x-5=6x-8\)

Mehek (mehek14):

now you can solve for x 6x - 3x = 3x 8 + -5 = 3 3 = 3x 1= x

Mehek (mehek14):

plug 1 for x in the first equation

Mehek (mehek14):

3*1 = 3 - 5 y = -2

Mehek (mehek14):

and when u graph them, the solution is (1,-2)

OpenStudy (anonymous):

Thank you

Mehek (mehek14):

your welcome

OpenStudy (anonymous):

What would part b be?

Mehek (mehek14):

at point (1,-2)

Mehek (mehek14):

because the solution is (1,-2)

OpenStudy (anonymous):

oh Thank you